# DEER: A Comprehensive and Reliable Benchmark for Deep-Research Expert Reports

### v1

@@ E.2 Coverage and Quality Factors
The rubric factors corresponding to Level 4 of the rubric evaluate each evaluation element (Level 3) from two perspectives: Coverage(C) and Quality(Q). The Coverage factor examines whether the element has been covered without omission as required throughout the document, i.e., how faithfully it includes content scattered in multiple locations or composed of multiple detailed requirements. On the other hand, the Quality factor evaluates whether the description possesses sufficient depth, logic, and rigor based on professional report standards, assuming the parts already written, i.e., "how well it is written." By separating Coverage and Quality in this way, we can independently measure (1) what is missing and (2) what the level of the included content is in a long report. Both factors use a 1–10 point scale, and detailed criteria for scoring are presented in Table 12.
@@ F.1 Overview
DREX’s Information Verification Protocol is based on a stepwise information verification procedure performed by human evaluators. This section details the actual two-step procedure (Claim Extraction, Factual Accuracy Evaluation) performed by human evaluators.
@@ F.3 Step 2: Claim Verification
Fact verification was strictly performed according to the following detailed criteria:
@@
  • Not Supported: When the basis for the claim cannot be found in the cited document, or the document is irrelevant to the topic. Cases where the document link is invalid (404) or the content cannot be verified are also included.
@@ G.2 Claim Extraction and Classification
To mitigate the "Lost-in-the-Middle" phenomenon that occurs when processing long contexts and to increase cost efficiency, this study introduced a Batch Extraction strategy. After dividing the entire report $D$ into sentence units $S=\{s_{1},s_{2},\dots,s_{m}\}$, they are processed in batches of a fixed size $B$ (in this study, $B=20$). Each batch processing provides the full report context ($D$) at the beginning of the prompt, but instructs the model to extract claims only for the sentences belonging to the current batch ($s_{i},\dots,s_{i+B-1}$).
@@
Classification Performance.
The results of performing type classification on 728 claims using GPT-5-mini are shown in Table 15. Explicit Citation (A) recorded a very high F1 score of 0.95, being identified almost perfectly. The distinction between Implicit Citation (BC) and Internal Information (DE) was somewhat challenging, but it recorded meaningful classification performance with F1 scores of 0.52 and 0.53, respectively. In particular, for the Unknown Source (F) type, which is the most difficult to identify, it achieved a precision of 0.35 and a recall of 0.25, indicating the ability to detect potential hallucination sentences. The overall accuracy was 0.68, which is a decent level considering the context of complex academic reports.
Table 15: Claim Classification Performance (GPT-5-mini). Classes are merged into A (Explicit), BC (Implicit), DE (Internal/Structural), and F (Unknown).
Class Precision Recall F1-Score Support
A (Explicit) 0.92 1.00 0.95 272
BC (Implicit) 0.42 0.68 0.52 163
DE (Internal) 0.76 0.41 0.53 242
F (Unknown) 0.35 0.25 0.30 51
Overall 0.68 0.68 0.67 728
@@
Table 16: Baseline comparison on correctly predicted B/C claims ($N=131$). The GPT-5-mini method achieves the best balance of precision and accuracy (Jaccard).
@@ G.3 Claim Verification
Relevant Context Filtering (Retrieval-Augmented Verification)
Using the entire report or long retrieved documents as input in the verification step is costly and can induce hallucinations due to unnecessary information (Noise). To solve this, we apply Relevant Context Filtering. For the verification target claim $q$ and the retrieved document $D_{retrieved}$, the document is divided into chunks $K=\{k_{1},k_{2},\dots,k_{r}\}$ (Size $\approx$ 1000 tokens). Then, the relevance score $Sim(q,k_{j})$ between each chunk and the claim is calculated, and only the top $N$ (Top-K) chunks are selected and used as input for the verification model.
In this study, the BM25 robertson2009probabilistic algorithm was used as the default Selector: The selected chunk set $K_{selected}=\{k\in K\mid rank(Sim(q,k))\leq N\}$ is combined while maintaining the original document order to form the final context $C_{final}$.
@@
Fair Use and Direct Quote Detection
To ensure compliance with copyright and research ethics, the appropriateness of direct quotes is evaluated. First, the similarity between sentences in the report and the original text (Reference) is calculated using RapidFuzz and SequenceMatcher. If the similarity ($\tau$) is 0.9 or higher, it is considered a direct quote. Sentences detected as direct quotes are primarily checked for quotation marks (" ") or block quote formatting using regular expressions. If the format is insufficient or additional judgment is needed, the LLM is presented with the sentence and the original text to comprehensively evaluate compliance with "Fair Use Guidelines" (citation indication, minimal quotation, etc.). The final score is calculated by deducting points from a maximum of 10 based on the non-compliant ratio among all sentences.
@@ G.4 Evaluation Metrics
  • External Claim Accuracy: The proportion of claims verified as factual among claims requiring external evidence (Type A, B, C).
@@
  • Citation Accuracy: The proportion of citations that correctly support the corresponding claim among all citations.
    $\text{Cit. Acc}=\frac{|\text{Supported Citations}|}{|\text{Total Citations}|}$
  • Reference Accuracy (Supported per Shown): The proportion of references that actually contributed to content verification (Supported) among unique references shown in the report.
    $\text{Ref. Acc}=\frac{|\text{Supported Unique References}|}{|\text{Total Unique References Shown}|}$
@@
    $\text{Reproducibility}=1-\frac{|\text{Error References}|}{|\text{Used References}|}$
  • Information Reliability: The proportion of references that are both reliable sources and support the content among the used references.
    $\text{Reliability}=\frac{|\text{Reliable \& Supported References}|}{|\text{Used References}|}$
  • Reference Diversity (Citations CV): Uses the Coefficient of Variation of citation counts to measure dependency on specific references. (Lower is better; inverted for scoring)
    $\text{CV}=\frac{\sigma(\text{Citation Counts})}{\mu(\text{Citation Counts})}$
@@
  • Externally Verifiable Claims Ratio: The proportion of claims verifiable with external evidence (Type A, B, C) among all claims.
    $\text{Verifiable Ratio}=\frac{|\text{Claims (A, B, C)}|}{|\text{Total Claims}|}$
@@
  • Information Quantity: The total number of claims verified as factual (Supported). (Raw Count before normalization)
    $\text{Info. Qty}=|\text{Accurate Verifiable Claims}|$
@@
  • Citation Quantity: The total number of valid citations (Supported Citation) supporting claims.
    $\text{Cit. Qty}=|\text{Supported Citations}|$
@@
  • Reference Quantity: The total number of valid references (Supported Reference) supporting claims.
    $\text{Ref. Qty}=|\text{Supported References}|$

### v2

@@ E.2 Coverage and Quality Factors
The rubric factors corresponding to Level 4 of the rubric evaluate each evaluation element (Level 3) from two perspectives: Coverage (C) and Quality (Q). The Coverage factor examines whether the element has been covered without omission as required throughout the document, i.e., how faithfully it includes content scattered in multiple locations or composed of multiple detailed requirements. On the other hand, the Quality factor evaluates whether the description possesses sufficient depth, logic, and rigor based on professional report standards, assuming the parts already written, i.e., “how well it is written.” By separating Coverage and Quality in this way, we can independently measure (1) what is missing and (2) what the level of the included content is in a long report. In our rubric, Level 4 consists of 101 factor-level items in total, comprising 66 Coverage items and 35 Quality items. Both factors use a 1–10 point scale, and sub-dimensions for scoring are presented in Table 12.
@@ F.1 Overview
DEER’s Information Verification Protocol is based on a stepwise information verification procedure performed by human evaluators. This section details the actual two-step procedure (Claim Extraction, Factual Accuracy Evaluation) performed by human evaluators.
@@ F.3 Step 2: Claim Verification
Fact verification was strictly performed according to the following sub-dimensions:
@@
  • Not Supported: When the basis for the claim cannot be found in the cited document, or the document is irrelevant to the topic.
  • Error: When the verification process fails due to accessibility issues (e.g., HTTP 4xx/5xx errors, Paywall, Captcha) or processing errors, preventing content verification.
Source Reliability Check
Independently of the content verification, we also evaluate the trustworthiness of the source domain itself.
  • Reliable: Trustworthy sources such as academic journals, official statistics, and authoritative institutions.
  • Unreliable: Sources with low credibility, such as personal blogs, social media posts, or unverified community forums.
@@ G.2 Claim Extraction and Classification
To mitigate the “Lost-in-the-Middle” phenomenon that occurs when processing long contexts and to increase cost efficiency, this study introduced a Batch Extraction strategy. After dividing the entire report $D$ into sentence units $S=\{s_{1},s_{2},\dots,s_{m}\}$, they are processed in batches of a fixed size $B$ (in this study, $B=20$). Each batch processing provides the full report context ($D$) at the beginning of the prompt, but instructs the model to extract claims only for the sentences belonging to the current batch ($s_{i},\dots,s_{i+B-1}$).
@@
Claim Extraction Evaluation Setup
To rigorously evaluate the claim extraction performance, we employed an LLM-based Judge (GPT-5). Standard lexical metrics (e.g., ROUGE Lin (2004), Exact Match) are unsuitable for this task because the generated claims may differ in wording or granularity (e.g., one sentence split into multiple atomic claims) while preserving the same semantic meaning. We defined two key metrics: (1) Semantic Recall: Measures whether each ground truth claim is semantically covered by the set of extracted claims. The LLM Judge compares the ground truth claim against all candidate claims extracted from the same source sentence and determines if the core information is present. (2) Classification F1: Measures whether the LLM correctly classified the claim type for the extracted claims. The implementation code and LLM prompts used for the evaluation are included in the supplementary material.
@@
Table 15: Baseline comparison on correctly predicted B/C claims ($N=131$). Backtracking achieves the best balance of precision, recall, and Jaccard.
@@ G.3 Claim Verification
Context Retrieval
Using the entire report or long retrieved documents as input in the verification step is costly and can induce hallucinations due to unnecessary information (Noise). To solve this, we apply Context Retrieval. For the verification target claim $q$ and the retrieved document $D_{retrieved}$, the document is divided into chunks $K=\{k_{1},k_{2},\dots,k_{r}\}$ (Size $\approx$ 1000 tokens). Then, the relevance score $Sim(q,k_{j})$ between each chunk and the claim is calculated, and only the top $N$ (Top-K) chunks are selected and used as input for the verification model.
In this study, the BM25 Robertson et al. (2009) and OpenAI’s text-embedding-3-large OpenAI was used as the embedding model: The selected chunk set $K_{selected}=\{k\in K\mid rank(Sim(q,k))\leq N\}$ is combined while maintaining the original document order to form the final context $C_{final}$.
@@
Augmented Dataset and Robustness Evaluation
While the human-annotated dataset provides high-quality ground truth, it exhibits significant class imbalance, with 82 “Supported” claims and only 5 “Not Supported” claims among 100 examples. This skew limits the ability to effectively evaluate the model’s capacity to discern unsupported claims and mitigate hallucinations. To address this, we constructed an adversarial augmented dataset. This dataset was generated by systematically perturbing initially supported claims—specifically by negating semantic meanings or altering numerical values—to create plausible but factually incorrect statements (i.e., “Not Supported”). All augmented examples were rigorously reviewed by human evaluators to ensure they are strictly false or unsupported by the source text.
Ablation Study on Retrieval Parameters
To identify the optimal configuration for the cost-efficient gpt-5-mini model, we conducted an ablation study varying Batch Size (10, 20), Reasoning Effort (Low, Medium, High), Retrieval Method (BM25 Robertson et al. (2009), OpenAI’s text-embedding-3-large OpenAI ), and Context Size (Top-K=2, 4). Table 16 summarizes the results on both the Original and Adversarial (Augmented) datasets. We observed that increasing the retrieved context size from Top-K=2 to 4 improved accuracy on the Original dataset (e.g., 77.0% $\rightarrow$ 79.3% for OpenAI, Low Effort), but increased the cost by approximately 35% (\$0.95 $\rightarrow$ \$1.28). The Low Reasoning Effort setting proved to be highly cost-effective, achieving comparable or superior performance to Medium/High effort while costing significantly less. Notably, the model demonstrated high robustness on the Adversarial dataset, maintaining high accuracy (>>88%) across most configurations, suggesting that it effectively distinguishes unsupported claims even under perturbations. Consequently, prioritizing feasibility of large-scale verification (cost/throughput) while retaining strong robustness, we selected the configuration Batch 20, Low Effort, Top-K=2, and OpenAI Embedding. This setup offers a minimal cost of \$0.95 per 1k claims while maintaining a strong accuracy of 77.0% (Original) and 88.5% (Adversarial), making it the most balanced choice for our resource-constrained high-volume verification pipeline.
Original Adversarial
Batch Effort Embed. Top-K Cost/1k (\$) Acc (%) F1 Acc (%) F1
10 high BM25 4 3.61 77.01 87.25 87.36 85.16
10 high OpenAI 4 3.67 78.16 88.00 88.46 86.62
10 low BM25 4 0.91 79.31 88.00 87.36 85.16
10 low OpenAI 4 1.03 78.16 88.00 89.01 87.34
10 medium BM25 4 1.75 73.56 84.93 87.36 85.16
10 medium OpenAI 4 1.62 73.56 84.93 87.91 85.90
20 high BM25 2 3.08 79.31 88.16 88.46 86.79
20 high BM25 4 3.38 79.31 88.74 89.56 88.05
20 high OpenAI 2 3.34 77.01 87.25 85.71 83.33
20 high OpenAI 4 3.31 75.86 86.49 88.46 86.79
20 low BM25 2 1.02 73.56 84.93 90.11 88.61
20 low BM25 4 1.30 81.61 90.20 89.01 87.34
20 low OpenAI 2 0.95 77.01 87.25 88.46 86.79
20 low OpenAI 4 1.28 79.31 88.74 89.56 87.90
20 medium BM25 2 1.57 74.71 85.71 86.81 84.42
20 medium BM25 4 1.79 79.31 88.74 88.46 86.62
20 medium OpenAI 2 1.43 73.56 84.93 87.91 86.08
20 medium OpenAI 4 1.71 74.71 85.71 88.46 86.79
Table 16: Ablation study of GPT-5-mini on Context Retrieval setting. Showing the impact of Batch Size, Reasoning Effort, Embedding Method, and Top-K chunks on performance. Original refers to the standard dataset, and Adversarial refers to the augmented dataset.
@@ G.4 Evaluation Metrics
  • Claim Factuality: The proportion of claims verified as factual among claims requiring external evidence (Type A, B, C).
@@
  • Citation Support: The proportion of citations that correctly support the corresponding claim among all citations.
    $\text{Score}=\frac{|\text{Supported Citations}|}{|\text{Total Citations}|}$
  • Reference Support: The proportion of references that actually contributed to content verification (Supported) among unique references shown in the report.
    $\text{Score}=\frac{|\text{Supported Unique References}|}{|\text{Total Unique References Shown}|}$
@@
    $\text{Score}=1-\frac{|\text{Error References}|}{|\text{Used References}|}$
  • Reference Reliability: The proportion of references that are both reliable sources and support the content among the used references.
    $\text{Score}=\frac{|\text{Reliable \& Supported References}|}{|\text{Used References}|}$
  • Reference Diversity (Normalized HHI): Measures how evenly citations are distributed across used references using the Normalized Herfindahl-Hirschman Index Rhoades (1993).
    $\text{Score}=10\times\left(1-\frac{\text{HHI}-1/N}{1-1/N}\right)$
@@
  • Evidence Coverage: The proportion of claims verifiable with external evidence (Type A, B, C) among all claims.
    $\text{Score}=\frac{|\text{Claims (A, B, C)}|}{|\text{Total Claims}|}$
@@
  • Information Amount: The total number of claims verified as factual (Supported).
    $\text{Score}=|\text{Accurate Verifiable Claims}|$
@@
  • Citation Amount: The total number of valid citations (Supported Citation) supporting claims.
    $\text{Score}=|\text{Supported Citations}|$
@@
  • Reference Amount: The total number of valid references (Supported Reference) supporting claims.
    $\text{Score}=|\text{Supported References}|$
Final Score Calculation
The final scores for Integrity and Sufficiency are computed by hierarchically aggregating the metrics (Metric $\rightarrow$ Criterion $\rightarrow$ Dimension), consistent with the score_avgs and criteria_avgs structure in the output.
1. Normalization (Metric Level)
Each raw metric value is first converted to a 0–10 scale:
  • Ratio-based metrics (e.g., Factuality): Scaled linearly.
    $\text{Score}=\min(\max(R,0),1)\times 10$
  • Quantity-based metrics (e.g., Counts): Scored via step function with divisors $D$ (Info=15, Cit=10, Ref=4).
    $\text{Score}=\min\left(\left\lfloor\frac{\max(N-1,0)}{D}\right\rfloor+1,10\right)$
2. Aggregation
  • Criterion Level: Average of normalized metric scores within each criterion.
  • Dimension Level: Average of criterion scores within each dimension.
Appendix H Baseline Model Details
We use the following backbone model families in our experiments: Qwen3-235B, Gemini 2.5, Claude Opus 4.5, and GPT-5. For readability, we refer to the GPT family as GPT-5 in the paper, while the actual backbone used in our runs was GPT-5.2. Since Gemini 2.5 Pro includes reasoning by default, we use Gemini 2.5 Flash for the fast (non-reasoning) setting, and Gemini 2.5 Pro for the other settings. For think, we use each service’s default reasoning budget. For think+search, we do not build a custom retrieval pipeline; instead, we use the built-in web search system provided by each service. We exclude Gemini think+search because citation information is not provided in its outputs. For WebThinker, we use Qwen3-235B as an auxiliary model. For OpenAI Deep Research, we collect reports generated from the service environment as of August 2025.
Appendix I Human-Correlation Experiment Setup
We measure alignment between LLM judges and expert human judgments on 45 reports. We consider five domains and sample three tasks from each, yielding 15 tasks in total. For each task, we randomly sample three reports from those produced by five Deep Research systems—OpenAI Deep Research, Gemini 2.5 Pro Deep Research, Claude Opus 4.1 Deep Research, WebThinker, and Qwen3-235B Deep Research—resulting in $15\times 3=45$ reports overall. WebThinker Li et al. (2025b) uses Qwen3-235B as an auxiliary model.
Each report was independently evaluated by two domain experts matched to the report’s topic, supporting both LLM–human correlation and human–human reliability analyses (90 ratings total; 45 reports × 2 experts). Experts had at least a master’s degree in a relevant field or comparable professional experience and, taking various factors into account, assigned an overall report quality score on a 1–5 scale (fractional values allowed). We used the mean of the two expert ratings as the report-level human score for LLM–human correlation, and we also reported agreement between the two experts’ ratings. Evaluating a report took 1.5 hours on average. Compensation followed the vendor’s standard payment framework, and we verified adherence to relevant procedures and policies.
We measure correlation with human judgments for five evaluator models (GPT-5, GPT-5-mini, Claude Opus 4.1, Claude Sonnet 4.5, and Gemini 2.5 Pro). Each model evaluates the same set of reports under each evaluation setting (Vanilla, +Dimensions, +Granular Rubrics, +Expert Guidance). Under the +Dimensions setting, evaluation is performed at the level of five report-quality dimensions. Under +Granular Rubrics and +Expert Guidance, evaluation is performed down to the level of rubric items that further break down those dimensions.
To measure alignment with expert human judgments, we report Pearson correlation ($r$), Spearman rank correlation ($\rho$), and pairwise agreement (PA). Each task corresponds to a single query and contains three reports. For each task, we compute Pearson $r$ and Spearman $\rho$ between human and model scores across the three reports, and report the mean over the 15 tasks. Following DeepBench Du et al. (2025), PA is the fraction of report pairs within a task for which the model’s relative preference matches the human relative preference; with three reports, each task has $\binom{3}{2}=3$ pairs. We compute PA per task and report the mean across tasks. To match the human 1–5 scale, we rescale model scores to 1–5 by dividing 1–10 scores by 2. For PA, a pair is counted as an agreement if the model and human judgments induce the same relation (>>, <<, or ==) for that pair. Tasks with undefined correlations (e.g., zero variance) are excluded from the corresponding averages.
For each setting, we compute the three metrics ($r$, $\rho$, and PA) independently for each of the five evaluator models, and then report the average across the models.
Appendix J Ethical Considerations and Potential Risks
DEER is designed as a research evaluation benchmark rather than a deployment or decision-making system, and its scores should be interpreted accordingly. Because the framework partially relies on LLM-based judges, evaluation outcomes may reflect systematic biases or limitations of the underlying models; however, this risk is mitigated through expert-authored rubrics, task-specific expert guidance, and validation against human expert judgments. Unless otherwise specified, reported results correspond to single-run evaluations under fixed configurations, rather than aggregates over multiple random seeds. DEER evaluates the quality and evidence alignment of generated reports, not the real-world correctness, safety, or deployability of the underlying claims or recommendations, and thus should not be used as an endorsement for practical application in high-stakes domains. In addition, while the information verification module checks both explicit and implicit claims against external sources, its reliability is inherently bounded by the quality, availability, and temporal stability of retrieved web evidence. Overall, DEER is intended for comparative research analysis and diagnostic evaluation, and any use beyond this scope should be accompanied by appropriate human oversight and domain-specific safeguards.
Appendix K Information About Use of AI Assistants
AI assistants were used in a limited and supportive manner during the preparation of this work. Specifically, they were employed to assist with coding-related tasks (e.g., drafting boilerplate scripts, refactoring, and debugging) and with writing-related support, such as grammar checking, stylistic refinement, and clarity improvement. Importantly, all conceptual contributions, experimental design, analysis, interpretation of results, and final manuscript were carried out by the authors.